Circular No. 5984

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET MAURY (1994h) This comet has been recovered by J. Scotti with the Spacewatch tele-

scope at Kitt Peak, the comet showing a 9′′ coma and a 0′.25 tail in p.a. 259^{o}. The following positions indicate a correction of ΔT = −0.52 day to the prediction on MPC 18257 (ephemeris on MPC 23142):

1994 UT  α_{2000}  δ_{2000}  m_{1}
May 3.45755  22^{h}42^{m}42.^{s}23  −0^{o}54′19′′.6
3.46074  22 42 42.64  −0 54 17.7  17.9
4.46252  22 44 37.28  −0 44 29.1  17.7
4.46935  22 44 38.04  −0 44 24.9

SUPERNOVA 1994M IN NGC 4493 B. Schmidt and R. Kirshner, Harvard-Smithsonian Center for Astro-

physics (CfA), report that a spectrum (range 370–760 nm) of SN 1994M taken by J. Peters (CfA) with the Tillinghast 1.5-m reflector (+ FAST spectrograph) on May 4.14 UT shows this object to be a type-Ia supernova, very close to maximum light. SN 1994M is located at α = 12^{h}28^{m}34^{s}.96, δ = +0^{o}52′53′′.4 (equinox 1950.0), as measured from a V CCD image taken by I. Dell’Antonio (CfA) with the F. L. Whipple Observatory’s 1.2-m telescope.

PERIODIC COMET KUSHIDA (1994a)

Ephemeris from orbital elements on MPC 23322:

1994TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
May 8 10^{h}29.^{m}50 + 4^{o}01.^{′}1 1.524 2.099 110.^{o}2 26.^{o}8 14.8
13 10 35.66 + 3 41.3 1.610 2.136 107.0 26.9 15.0
18 10 41.99 + 3 19.0 1.697 2.172 103.8 26.9 15.2
23 10 48.44 + 2 54.4 1.785 2.209 100.6 26.8 15.4
28 10 55.00 + 2 27.6 1.876 2.246  97.5 26.6 15.6
June 2 11 01.65 + 1 59.1 1.967 2.282  94.5 26.3 15.8
7 11 08.37 + 1 28.8 2.060 2.319  91.4 25.9 16.0
12 11 15.15 + 0 57.0 2.154 2.356  88.4 25.5 16.1
17 11 21.99 + 0 23.8 2.248 2.392  85.4 25.0 16.3
22 11 28.86 − 0 10.5 2.343 2.429  82.5 24.5 16.5
27 11 35.77 − 0 45.9 2.438 2.465  79.5 23.9 16.6

1994 May 4 Daniel W. E. Green